OpenStudy (anonymous):

Solve the equation by completing the square: x2 + 2x = 120 Choose one answer. a. x = -12 or 10 b. x =-12 or -10 c. x =12 or x =10 d. x = 12 or -10

OpenStudy (anonymous):

\[x^2+2x=120\] \[(x+1)^2=120+1=121\] \[x+1=\pm\sqrt{121}\] \[x+1=11\] or \[x+1=-11\] \[x=10\] or \[x=-12\]

OpenStudy (anonymous):

oh thats wat i did wrong i got 10 or 12 for my answer thnk u